OpenStudy (anonymous):

Factor x^2-12x+36-49y^2

OpenStudy (anonymous):

I have x^2+24x-49y^2 so far

OpenStudy (anonymous):

\[a^2-b^2=\left( a+b \right)\left( a-b \right)\] \[x^2-12x+36-49y^2=\left( x-6 \right)^2-\left( 7y \right)^2=?\]

OpenStudy (anonymous):

um x^2-12x-36-49y^2

OpenStudy (anonymous):

I'm lost

jimthompson5910 (jim_thompson5910):

Hint: \[\Large (x-6)^2 - (7y)^2\] \[\Large ({\color{red}{x-6}})^2 - ({\color{blue}{7y}})^2\] \[\Large ({\color{red}{a}})^2 - ({\color{blue}{b}})^2 = ({\color{red}{a}}-{\color{blue}{b}})({\color{red}{a}}+{\color{blue}{b}})\]

jimthompson5910 (jim_thompson5910):

\[\Large x^2 - 12x + 36 = (x-6)^2\] \[\Large 49y^2 = (7y)^2\]

OpenStudy (anonymous):

Would that be all?

jimthompson5910 (jim_thompson5910):

no there's one more step

OpenStudy (anonymous):

Hit me :)

jimthompson5910 (jim_thompson5910):

use the difference of squares rule to factor \[\Large (x-6)^2 - (7y)^2\]

OpenStudy (anonymous):

(x-6)^2=x^-12x-36 right? (7y)^2=49y^2 right?

jimthompson5910 (jim_thompson5910):

you're going backwards though

jimthompson5910 (jim_thompson5910):

what I'm aiming for is to go from this \[\Large ({\color{red}{x-6}})^2 - ({\color{blue}{7y}})^2\] to this \[\Large ({\color{red}{x-6}}-{\color{blue}{7y}})({\color{red}{x-6}}+{\color{blue}{7y}})\] I used the difference of squares rule

jimthompson5910 (jim_thompson5910):

it's like saying x^2 - 64 factors to (x-8)(x+8)

OpenStudy (anonymous):

Oh!! I get it now haha :D Thank you!

jimthompson5910 (jim_thompson5910):

np